OpenStudy (sydneeod):

What is the value of f ′(5) for f(x) = 3x^2 – 2x – 4? I don't want the answer, but just help with the steps of setting up the problem please

Nnesha (nnesha):

first take the derivative of the function and then substitute 5 for x

Nnesha (nnesha):

\[\huge\rm f(x) = \color{blue}{3x^2 – 2x – 4}\] what would be the derivative of this function ??

OpenStudy (sydneeod):

what equation do you use to find the derivative? f(x+h)-f(x)/h ?

OpenStudy (sydneeod):

@Nnesha

Nnesha (nnesha):

same equation for f(x+h) replace `x` with `x+h`

Nnesha (nnesha):

oh no the given function f(x) = 3x^2 – 2x – 4

OpenStudy (sydneeod):

so 3(5+h)^2-2(5+h)-4 -(3(5)^2-2(5)-4)/h ?

OpenStudy (sydneeod):

could I also use this formula? f(x)-f(a)/x-a ? or is that for something entirely different?

Nnesha (nnesha):

we have to find the value of f'(5) where x=5 so \[\huge\rm \frac{ f(\color{Red}{x}+h)-f(\color{Red}{x}) }{ h }\] \[\huge\rm \frac{ f(\color{Red}{5}+h)-f(\color{Red}{x}) }{ h }\] to find the value of f'(5) we have to take the limit as x approaches 0 \[\large\rm f'(5)=\lim_{h \rightarrow 0}\frac{ f(\color{Red}{5}+h)-f(\color{Red}{x}) }{ h }\]

Nnesha (nnesha):

correction *** \[\huge\rm \frac{ f(\color{Red}{5}+h)-f(\color{Red}{5}) }{ h }\] and i meant to say `h approaches 0`

OpenStudy (sydneeod):

ooooh okay, so find out what 3(5)^2-2(5)-4 =, 36, and then it looks like (36+h)-36 / h?

Nnesha (nnesha):

to find the derivative of a function at some point (f'(a)) first step) find `f(a)` 2nd step) find `f(a+h)` 3rd) use the formula \[\frac{ f(a+h)-f(a) }{ h }\] and then take the limit as h->0

Nnesha (nnesha):

hmm i didn't get 36..

OpenStudy (sydneeod):

oops, 61? must have multiplied wrong

Nnesha (nnesha):

yeah looks good now find f(a+h) for your question a=5 so substitute `5+h` into the original equation

Nnesha (nnesha):

\[\large\rm f'(5)= \lim_{h \rightarrow 0} \frac{ f(\color{blue}{5+h})- f(\color{Red}{a}) }{ h }\]\[\large\rm f'(5)= \lim_{h \rightarrow 0} \frac{[\color{blue}{3(5+h)^2-2(5+h)-4} ]- \color{Red}{61} }{ h }\] simplify :=))

Nnesha (nnesha):

confused ??? ~o.0~

Nnesha (nnesha):

tag me if u have question abt it :=))

OpenStudy (sydneeod):

I believe the answer is 28 @Nnesha very confused though haha but you've helped a lot!

Nnesha (nnesha):

confused about what ?

Nnesha (nnesha):

\[\large\rm \lim_{h\rightarrow 0 }\frac{ 3(25+10h+h^2)-10-2h -61}{ h}\]\[\large\rm \lim_{h\rightarrow 0 }\frac{ 75+30h+3h^2-10-2h -61}{ h}\] simplify the h canceled out (at the denominator ) then substitute h for 0 you will find the value of f'(5) :=)) good job!

OpenStudy (mathmale):

The approach you're using here is called "definition of the derivative." Were you told to use this method? If you have already learned some of the basic derivative rules, try tackling this problem with them. Pls. show your work.